selling price per unit 24

Variable expense per unit 15

fixed expense per month 7470

unit sales per month 980

1. What is the company’s margin of safety? (Do not round intermediate calculations.)

2. What is the company’s margin of safety as a percentage of its sales? (Round your percentage answer to 2 decimal places (i.e. 0.1234 should be entered as 12.34).)
Answer #1

Answers :

Margin of Safety Units : 150 Units

MArgin of Safty Dollars: $ 3,600

MArgin of Safety % = 15.31%

1) MArgin of safety = Actual Sales-BEP Sales

BEP sales = fixed Cost / CM per unit = 7470/(24-15) = 830 Units

Margin of Safety Units : 980-830 = 150 units

margin of Safety dollars: 150*24 = $3,600

MArgin of Safety % = 150/980 = 15.31%